(a) Calculate the magnitude of the gravitational force exerted on a 1,412-kg satellite that is a distance of two earth radii from the center of the earth.
________ N

(b) What is the magnitude of the gravitational force exerted on the earth by the satellite?
__________ N

(c) Determine the magnitude of the satellite's acceleration.
_______ m/s2

(d) What is the magnitude of the earth's acceleration?
_______ ms2
Answer #1

A)

Force, F = G m1 m2/r^2

F = (6.67 x 10^-11 x 5.98 x 10^24 x 1412)/(2 x 6.38 x 10^6)^2

F = 3459.08 N

B)

Force, F = 3459.08 N

C)

Acceleration, a = (3459.08)/1412 = 2.45 m/s^2

D)

Acceleration, a = (3459.08)/(5.98 x 10^24) = 5.78 x 10^-22 m/s^2

Comment in case any doubt please rate my answer ....